# RichControl: Structure- and Appearance-Rich Training-Free Spatial Control for Text-to-Image Generation

### v3

@@ Abstract
Figure 1: Structure- and appearance-rich training-free spatial control for text-to-image generation. We propose a training-free framework that enables high-quality spatial control for pretrained text-to-image diffusion models under arbitrary spatial conditions. (Left) By introducing strong and intuitive structure and appearance control, our method effectively addresses key limitations of prior work such as Ctrl-X [47], including structure misalignment, condition leakage, and artifacts, and (Right) achieves SOTA performance among all training-free methods; in the radar chart, greater distance from the center indicates superior results.
@@ 1 Introduction
With the success of text-to-image (T2I) diffusion models [70, 64, 61], recent research has explored integrating conditional images, e.g., depth maps for spatial control. Early approaches, such as ControlNet [102], rely on fine-tuning or auxiliary networks trained on paired data, which constrains their flexibility and scalability. More recent studies have shown that the rich structural information encoded within diffusion features can be exploited to guide image generation without retraining, thereby enabling zero-shot control. They either introduce additional guidance terms to minimize the feature distance between the condition and target during denoising [18, 6, 52], or inject features extracted from the condition image at each timestep into the target image [28, 80, 47]. Among them, feature injection-based methods such as Ctrl-X [47] have shown promising performance across diverse conditioning scenarios.
@@
However, these methods still encounter several failures, including structural misalignment, condition leakage, and visual artifacts (Fig. 1). These issues become more pronounced when the condition image deviates significantly from natural RGB distributions, e.g., in pose or depth maps (Fig. 5). This suggests that a key challenge lies in the domain gap between condition and natural image features in pretrained T2I diffusion models. We hypothesize that the injected condition features often lie outside the distribution of natural image features, which hinders the synthesis of high-fidelity results. This motivates us to analyze the temporal dynamics of diffusion features, observing a trade-off between structural fidelity and domain alignment (see Figs. 3 and 3). These findings expose a fundamental limitation in existing training-free methods [28, 80, 47], which rely on condition features extracted at the same timestep during denoising. This schedule fails to accommodate the evolving trade-off across timesteps: early features leads to loss of structural detail, while late features result in domain mismatch and condition leakage (Fig. 3).
@@
To address this, we generalize the sampling process of condition features and explore the design space of the feature injection schedule. The result shows that the optimal timestep is neither the same one as the target output image nor the latest one with the clearest features. Through a comprehensive investigation, we identify a family of candidate schedules that share an identical last timestep, among which a constant schedule yields consistently strong results. Building on these insights, we propose a more flexible feature injection framework that decouples the injection timestep from the denoising process. To further enhance control precision and visual fidelity, we apply a restart refinement schedule that iteratively mitigates visual artifacts introduced by injected features, and incorporate prompt augmentation to ensure semantic alignment with the condition image. Together, these designs enable structure- and appearance-richer control of pretrained diffusion models [61]. Fig. 4 provides an overview of our framework, which consists of three key components: (i) Structure-Rich Injection (SRI) injects condition features based on a principled sampling schedule; (ii) Restart Refinement (RR) performs iterative forward–backward denoising; (iii) Appearance-Rich Prompting (ARP) aligns the semantics of the appearance prompt with the condition image.
@@
Extensive experiments validate the effectiveness of our approach across diverse types of condition images, demonstrating improved structural consistency, visual fidelity, and semantic alignment compared to state-of-the-art training-free methods. Furthermore, the idea of our framework can be readily incorporated into other training-free methods, such as FreeControl [52], where it yields notable improvements, highlighting its versatility. In summary, our contributions are threefold: (i) We reveal the inherent limitation of existing training-free methods in sampling schedules, and identify a spectrum of alternatives through a principled analysis of the design space. (ii) Building on this insight, we propose a novel framework that enables structure- and appearance-rich controllable T2I generation. (iii) Our proposed method demonstrates state-of-the-art performance in comparison to previous training-based and training-free baselines, delivering superior structure preservation, text-image alignment, and visual fidelity.
@@ 2.1 T2I Diffusion Models
Text-to-image (T2I) diffusion models [31, 70, 63, 64] typically leverage U-Net [65] or transformer-based backbones [58, 19, 41] and integrate textual information via cross-attention or classifier-free guidance [64, 29, 54]. Latent diffusion models like Stable Diffusion [64] introduce compressed latent spaces to reduce computational cost. In addition to architectural innovations, some work focuses on improving sampling efficiency [72, 49, 36, 98, 48, 73, 109]. Restart Sampling [98] proposes alternating between adding noise and denoising to balance discretization error and contraction. Our work explores sampling strategies in the context of conditional text-to-image generation.
@@ 2.2 Training-based Controllable Diffusion Models
It is difficult to convey human preferences through text descriptions alone. Training-based controllable diffusion models mitigate this problem by training auxiliary modules or fine-tuning the model to incorporate additional input signals to guide the generation process. According to task characteristics, these methods can be broadly classified into three categories: (i) Image editing [7, 26, 37, 83, 71, 24, 92, 88, 12, 42, 90, 94, 27] takes an input image and applies targeted modifications while preserving other regions of the image; (ii) Image-to-image translation [34, 69, 79, 55, 56] learns mappings between images of different domains; (iii) Conditional text-to-image (T2I) generation methods synthesize images that satisfy both a text prompt and a control condition. Among these approaches, some works condition the generation on layout cues (e.g., bounding boxes) [46, 100, 84] or reference images of specific subjects [23, 67, 68, 3, 60, 44, 104, 107, 75, 76, 12, 90, 92, 88, 42]. Another line of work [102, 53, 101, 108, 4, 104, 107, 75, 76, 45, 92, 88, 12, 42, 90, 94, 27] enables fine-grained structural control by leveraging condition images of different modalities (e.g., canny edges, OpenPose keypoints [9]). Despite their impressive performance, these methods all require retraining or fine-tuning on datasets tailored to the control signal, which limits their generalization to new model checkpoints and novel control conditions.
@@ 2.3 Training-free Controllable Diffusion Models
On the other hand, training-free controllable diffusion models operate at inference time to achieve condition control without additional training on task-specific paired data. Similar to Sec. 2.2, we categorize them into 3 groups: (i) Image editing [8, 99, 18, 57, 105, 77, 35, 14, 21, 15, 111, 5, 82, 96, 97, 87, 78, 32]; (ii) Image-to-image translation [74], with some works [2, 47, 40, 33, 25, 13] further conditioning the transformation on an additional appearance image; (iii) Conditional text-to-image (T2I) generation, which generates images consistent with both textual prompts and input control signals. These signals range from coarse layout constraint such as bounding boxes [91, 11, 93, 43, 86], semantic references like subject images [106, 20, 85, 16, 59, 66], to condition images that provide fine-grained control [47, 52, 80, 28, 6, 51, 39]. These approaches enable flexible control over pre-trained diffusion models and can generalize to novel control modalities without the cost of additional data collection or retraining. As a training-free conditional T2I method conditioned on condition images, our approach extends this line of work by improving both control fidelity and generation quality across diverse visual conditions.
@@ 3 Revisiting Sampling Schedules of Condition Features
Given a text prompt $\mathcal{P}$ and a condition image $\mathbf{I}^{\text{struct}}$ of arbitrary modality, our goal is to generate an output image $\mathbf{I}$ that semantically aligns with the prompt $\mathcal{P}$ while preserving the structure of $\mathbf{I}^{\text{struct}}$. To align the structure of the generated image with that of the condition $\mathbf{I}^{\text{struct}}$, recent training-free approaches such as Ctrl-X [47] leverage the diffusion features of a noisy latent $\mathbf{x}_{t}^{\text{struct}}$ of the condition image. Specifically, they obtain a clean latent of the $\mathbf{x}_{0}^{\text{struct}}$ by encoding $\mathbf{I}^{\text{struct}}$ using a Variational Auto-Encoder, and then get its noisy version $\mathbf{x}_{t}^{\text{struct}}$ through DDIM inversion [72] or the diffusion forward process. Intermediate features are subsequently extracted from designated layers of the model backbone, and condition features are injected into those of $\mathbf{I}$ at each timestep. We denote these condition features as $\mathbf{f}_{l,t}^{\text{struct}}$, where $l$ refers to the layer index and $t$ denotes the timestep.
@@
While enabling zero-shot spatial control with diverse condition modalities, these methods often suffer from structural misalignment and condition leakage. For instance, Ctrl-X [47] fails to preserve the structure of the panda (Fig. 1). Empirically, we observe that these failures are further exacerbated when the condition image deviates substantially from natural RGB images, as in the case of pose or depth maps shown in Fig. 5. This suggests that a key challenge lies in the domain gap between the condition and natural image distributions in the feature space of pretrained diffusion models. We hypothesize that the injected features $\mathbf{f}_{l,t}^{\text{struct}}$ fall outside the distribution of natural image features, thereby reducing their effectiveness in preserving the structure of the condition image $\mathbf{I}^{\text{struct}}$ during generation.
@@ 4.1 Structure-Rich Injection
The structure-rich injection strategy adopts a sampling schedule where the extracted condition features are both semantically compatible and structurally informative. Specifically, we begin by encoding the structure condition image $\mathbf{I}^{\text{struct}}$ using the model backbone to obtain the condition features and attention maps, as illustrated in Fig. 4. Prior work [28, 80, 52, 47] typically extracts condition features $\mathbf{f}_{l,t}^{\text{struct}}$ and attention maps $\mathbf{A}_{l,t}^{\text{struct}}$ at the same timestep $t$ as the denoising step used for generating the output image $\mathbf{I}$ (the bottom branch of the U-Net in Fig. 4), where $l$ denotes the specific U-Net layer to be injected.
@@ 4.2 Restart Refinement
Despite a carefully designed injection schedule, structure-rich features can still introduce out-of-distribution artifacts and condition leakage during denoising. To address this, we adopt a restart refinement schedule inspired by diffusion-based sampling methods [98]. After several rounds of structure and appearance control, we inject noise at an intermediate timestep $t_{\text{min}}$, effectively restarting the denoising process by transitioning the latent to $t_{\text{max}}$ step. A DDIM backward step [72] is then applied. This forward–backward cycle is repeated $N$ times within $t_{\text{min}},t_{\text{max}}]$. In the $i^{\textrm{th}}$ iteration ($i\in\{0,\dots,N-1\}$), the restart proceeds as follows:
@@ 4.3 Appearance-Rich Prompting
To enhance the realism of the generated image, prior work [52, 47] generates an auxiliary appearance image $\mathbf{I}^{\text{app}}$ and perform appearance transfer (the middle branch of the U-Net in Fig. 4; see Appx. D for technical details). However, brief and ambiguous user prompts sometimes hinder the establishment of semantic correspondence between the appearance image $\mathbf{I}^{\text{app}}$ and the output image $\mathbf{I}$ in existing appearance transfer methods, leading to artifacts. For example, as illustrated in Fig. 4, the condition image is a frontal view of a cat’s head, while the text prompt specifies ”a bear”, resulting in an appearance image $\mathbf{I}^{\text{app}}$ that depicts a full-body bear.
To tackle this issue, we propose Appearance-Rich Prompting (ARP), a strategy that leverages multimodal large language models [1] to systematically align the semantics of $\mathcal{P}$ with those of the conditions $\mathbf{I}^{\text{struct}}$. Illustrative examples are provided in the ablation study in Fig. 8. Please refer to Appx. D for details of the prompt engineering pipeline $\mathcal{E}$.
@@ 5.1 Setup
We base our evaluation on datasets from prior work [52, 47]. However, many of the condition types in prior datasets are underrepresented, resulting in an imbalanced distribution. To enable more consistent evaluation, we collect seven commonly used structural conditions: canny edges, depth maps, normal maps, human poses, segmentation masks, HED edges, and scribble drawings, and construct a more balanced dataset with over 20 image-text pairs for each condition type.
@@
We evaluate our method against 6 existing training-free baselines: Ctrl-X [47], FreeControl [52], PnP [80], P2P [28], SDEdit [51], and InfEdit [97], as well as 2 training-based baselines: ControlNet [102] and T2I-Adapter [53]. Experiment results on four condition types supported by T2I-Adapter-SDXL (canny, depth, normal, and pose) are provided in Sec. F.3.
@@
Following prior work [52, 47], we employ three widely-adopted metrics for comparison. (i) CLIP score [62] measures text-image alignment via cosine similarity between the CLIP embeddings of the generated image and text prompt. (ii) DINO self-similarity score (Self-sim) [10, 79] quantifies structural alignment in the DINO-ViT feature space. (iii) Condition LPIPS score (LPIPS) [103] measures perceptual deviation between the generated image and the condition image.
In addition to traditional metrics, we further adopt three reward model metrics that more accurately reflect human preferences. (i) DreamSim [22] evaluates the perceptual similarity of two images by capturing both mid-level similarities (image layout, object pose, semantic content) and low-level attributes (color, texture). (ii) ImageReward [95] measures image quality and text-image alignment based on a reward model trained with human feedback. (iii) HPSv2 [89] serves as a reliable indicator of overall generation quality aligned with human judgments.

### v4

@@ Abstract
Figure 1: Structure- and appearance-rich training-free spatial control for text-to-image generation. We propose a training-free framework that enables high-quality spatial control for pretrained text-to-image diffusion models under arbitrary spatial conditions. (Left) By introducing strong and intuitive structure and appearance control, our method effectively addresses key limitations of prior work such as Ctrl-X [lin2024ctrlx], including structure misalignment, condition leakage, and artifacts, and (Right) achieves SOTA performance among all training-free methods; in the radar chart, greater distance from the center indicates superior results. Project page: https://zhang-liheng.github.io/richcontrol/.
@@ 1 Introduction
With the success of text-to-image (T2I) diffusion models [saharia2022imagen, rombach2022ldm, podell2023sdxl], recent research has explored integrating conditional images, e.g., depth maps for spatial control. Early approaches, such as ControlNet [zhang2023controlNet], rely on fine-tuning or auxiliary networks trained on paired data, which constrains their flexibility and scalability. More recent studies have shown that the rich structural information encoded within diffusion features can be exploited to guide image generation without retraining, thereby enabling zero-shot control. They either introduce additional guidance terms to minimize the feature distance between the condition and target during denoising [epstein2023selfguidance, bansal2023universalguidance, mo2023freecontrol], or inject features extracted from the condition image at each timestep into the target image [hertz2022prompt-to-prompt, tumanyan2023plug-and-play, lin2024ctrlx]. Among them, feature injection-based methods such as Ctrl-X [lin2024ctrlx] have shown promising performance across diverse conditioning scenarios.
@@
However, these methods still encounter several failures, including structural misalignment, condition leakage, and visual artifacts (Fig. 1). These issues become more pronounced when the condition image deviates significantly from natural RGB distributions, e.g., in pose or depth maps (Fig. 5). This suggests that a key challenge lies in the domain gap between condition and natural image features in pretrained T2I diffusion models. We hypothesize that the injected condition features often lie outside the distribution of natural image features, which hinders the synthesis of high-fidelity results. This motivates us to analyze the temporal dynamics of diffusion features, observing a trade-off between structural fidelity and domain alignment (see Figs. 3 and 3). These findings expose a fundamental limitation in existing training-free methods [hertz2022prompt-to-prompt, tumanyan2023plug-and-play, lin2024ctrlx], which rely on condition features extracted at the same timestep during denoising. This schedule fails to accommodate the evolving trade-off across timesteps: early features leads to loss of structural detail, while late features result in domain mismatch and condition leakage (Fig. 3).
@@
To address this, we generalize the sampling process of condition features and explore the design space of the feature injection schedule. The result shows that the optimal timestep is neither the same one as the target output image nor the latest one with the clearest features. Through a comprehensive investigation, we identify a family of candidate schedules that share an identical last timestep, among which a constant schedule yields consistently strong results. Building on these insights, we propose a more flexible feature injection framework that decouples the injection timestep from the denoising process. To further enhance control precision and visual fidelity, we apply a restart refinement schedule that iteratively mitigates visual artifacts introduced by injected features, and incorporate prompt augmentation to ensure semantic alignment with the condition image. Together, these designs enable structure- and appearance-richer control of pretrained diffusion models [podell2023sdxl]. Fig. 4 provides an overview of our framework, which consists of three key components: (i) Structure-Rich Injection (SRI) injects condition features based on a principled sampling schedule; (ii) Restart Refinement (RR) performs iterative forward–backward denoising; (iii) Appearance-Rich Prompting (ARP) aligns the semantics of the appearance prompt with the condition image.
@@
Extensive experiments validate the effectiveness of our approach across diverse types of condition images, demonstrating improved structural consistency, visual fidelity, and semantic alignment compared to state-of-the-art training-free methods. Furthermore, the idea of our framework can be readily incorporated into other training-free methods, such as FreeControl [mo2023freecontrol], where it yields notable improvements, highlighting its versatility. In summary, our contributions are threefold: (i) We reveal the inherent limitation of existing training-free methods in sampling schedules, and identify a spectrum of alternatives through a principled analysis of the design space. (ii) Building on this insight, we propose a novel framework that enables structure- and appearance-rich controllable T2I generation. (iii) Our proposed method demonstrates state-of-the-art performance in comparison to previous training-based and training-free baselines, delivering superior structure preservation, text-image alignment, and visual fidelity.
@@ 2.1 T2I Diffusion Models
Text-to-image (T2I) diffusion models [ho2022cascaded, saharia2022imagen, ramesh2022dalle2, rombach2022ldm] typically leverage U-Net [olaf2015unet] or transformer-based backbones [Peebles2023DiT, esser2024sd3, flux2024] and integrate textual information via cross-attention or classifier-free guidance [rombach2022ldm, ho2022cfg, nichol2021glide]. Latent diffusion models like Stable Diffusion [rombach2022ldm] introduce compressed latent spaces to reduce computational cost. In addition to architectural innovations, some work focuses on improving sampling efficiency [song2020ddim, lu2022dpmsolver, karras2022edm, xu2023restart, liu2023rectifiedflow, song2023consistency, zhao2023unipc]. Restart Sampling [xu2023restart] proposes alternating between adding noise and denoising to balance discretization error and contraction. Our work explores sampling strategies in the context of conditional text-to-image generation.
@@ 2.2 Training-based Controllable Diffusion Models
It is difficult to convey human preferences through text descriptions alone. Training-based controllable diffusion models mitigate this problem by training auxiliary modules or fine-tuning the model to incorporate additional input signals to guide the generation process. According to task characteristics, these methods can be broadly classified into three categories: (i) Image editing [brooks2023instructpix2pix, goel2023pairdiffusion, kim2022diffusionclip, wang2023imageneditor, sheynin2024emuedit, geng2024instructdiffusion, Xiao2025Omnigen, wu2025omnigen2, chen2025unireal, Le2025Onediffusion, Xia2025Dreamomni, xie2025anyrefill, han2024ace] takes an input image and applies targeted modifications while preserving other regions of the image; (ii) Image-to-image translation [isola2017im2im, saharia2022palette, tumanyan2022splicing, ouyang2025klora, park2019spade] learns mappings between images of different domains; (iii) Conditional text-to-image (T2I) generation methods synthesize images that satisfy both a text prompt and a control condition. Among these approaches, some works condition the generation on layout cues (e.g., bounding boxes) [li2023gligen, yang2023reco, wang2024instancediffusion] or reference images of specific subjects [gal2022ti, ruiz2023dreambooth, ruiz2023hyperdreambooth, avrahami2023break-a-scene, po2023orthogonal, li2023blipdiffusion, Zhang2024SSRencoder, zhang2025easycontrol, tan2025ominicontrol, tan2025ominicontrol2, chen2025unireal, Xia2025Dreamomni, Xiao2025Omnigen, wu2025omnigen2, Le2025Onediffusion]. Another line of work [zhang2023controlNet, mou2023t2i, ye2023ipadapter, zhao2023unicontrol, avrahami2023spatext, Zhang2024SSRencoder, zhang2025easycontrol, tan2025ominicontrol, tan2025ominicontrol2, li2024controlnet_plus_plus, Xiao2025Omnigen, wu2025omnigen2, chen2025unireal, Le2025Onediffusion, Xia2025Dreamomni, xie2025anyrefill, han2024ace] enables fine-grained structural control by leveraging condition images of different modalities (e.g., canny edges, OpenPose keypoints [cao2019openpose]). Despite their impressive performance, these methods all require retraining or fine-tuning on datasets tailored to the control signal, which limits their generalization to new model checkpoints and novel control conditions.
@@ 2.3 Training-free Controllable Diffusion Models
On the other hand, training-free controllable diffusion models operate at inference time to achieve condition control without additional training on task-specific paired data. Similar to Sec. 2.2, we categorize them into 3 groups: (i) Image editing [cao2023masactrl, xu2024headrouter, epstein2023selfguidance, parmar2023pix2pix-zero, zhang2024samcontrol, tewel2025addit, jia2024designedit, couairon2022diffedit, Feng2025Dit4edit, dalva2024fluxspace, zhu2025kvedit, Avrahami2025StableFlow, wang2024rfedit, Xu2025unveil, xu2023infedit, wei2025freeflux, titov2024guideandrescale, hu2025islock]; (ii) Image-to-image translation [su2022ddib], with some works [alaluf2023crossimage, lin2024ctrlx, kwon2023diffuseit, huang2025attenst, go2024eyeforaneye, chung2024styleid] further conditioning the transformation on an additional appearance image; (iii) Conditional text-to-image (T2I) generation, which generates images consistent with both textual prompts and input control signals. These signals range from coarse layout constraint such as bounding boxes [xiao2024rb, chen2023layoutguidance, xie2023boxdiff, li2025winwinlay, wang2025DADG], semantic references like subject images [zhang2025freecus, feng2025personalizeanything, wang2025freeevent, ding2024freecustom, pham2024tale, rout2025rbmodulation], to condition images that provide fine-grained control [lin2024ctrlx, mo2023freecontrol, tumanyan2023plug-and-play, hertz2022prompt-to-prompt, bansal2023universalguidance, meng2021sdedit, kim2023densediffusion]. These approaches enable flexible control over pre-trained diffusion models and can generalize to novel control modalities without the cost of additional data collection or retraining. As a training-free conditional T2I method conditioned on condition images, our approach extends this line of work by improving both control fidelity and generation quality across diverse visual conditions.
@@ 3 Revisiting Sampling Schedules of Condition Features
Given a text prompt $\mathcal{P}$ and a condition image $\mathbf{I}^{\text{struct}}$ of arbitrary modality, our goal is to generate an output image $\mathbf{I}$ that semantically aligns with the prompt $\mathcal{P}$ while preserving the structure of $\mathbf{I}^{\text{struct}}$. To align the structure of the generated image with that of the condition $\mathbf{I}^{\text{struct}}$, recent training-free approaches such as Ctrl-X [lin2024ctrlx] leverage the diffusion features of a noisy latent $\mathbf{x}_{t}^{\text{struct}}$ of the condition image. Specifically, they obtain a clean latent of the $\mathbf{x}_{0}^{\text{struct}}$ by encoding $\mathbf{I}^{\text{struct}}$ using a Variational Auto-Encoder, and then get its noisy version $\mathbf{x}_{t}^{\text{struct}}$ through DDIM inversion [song2020ddim] or the diffusion forward process. Intermediate features are subsequently extracted from designated layers of the model backbone, and condition features are injected into those of $\mathbf{I}$ at each timestep. We denote these condition features as $\mathbf{f}_{l,t}^{\text{struct}}$, where $l$ refers to the layer index and $t$ denotes the timestep.
@@
While enabling zero-shot spatial control with diverse condition modalities, these methods often suffer from structural misalignment and condition leakage. For instance, Ctrl-X [lin2024ctrlx] fails to preserve the structure of the panda (Fig. 1). Empirically, we observe that these failures are further exacerbated when the condition image deviates substantially from natural RGB images, as in the case of pose or depth maps shown in Fig. 5. This suggests that a key challenge lies in the domain gap between the condition and natural image distributions in the feature space of pretrained diffusion models. We hypothesize that the injected features $\mathbf{f}_{l,t}^{\text{struct}}$ fall outside the distribution of natural image features, thereby reducing their effectiveness in preserving the structure of the condition image $\mathbf{I}^{\text{struct}}$ during generation.
@@ 4.1 Structure-Rich Injection
The structure-rich injection strategy adopts a sampling schedule where the extracted condition features are both semantically compatible and structurally informative. Specifically, we begin by encoding the structure condition image $\mathbf{I}^{\text{struct}}$ using the model backbone to obtain the condition features and attention maps, as illustrated in Fig. 4. Prior work [hertz2022prompt-to-prompt, tumanyan2023plug-and-play, mo2023freecontrol, lin2024ctrlx] typically extracts condition features $\mathbf{f}_{l,t}^{\text{struct}}$ and attention maps $\mathbf{A}_{l,t}^{\text{struct}}$ at the same timestep $t$ as the denoising step used for generating the output image $\mathbf{I}$ (the bottom branch of the U-Net in Fig. 4), where $l$ denotes the specific U-Net layer to be injected.
@@ 4.2 Restart Refinement
Despite a carefully designed injection schedule, structure-rich features can still introduce out-of-distribution artifacts and condition leakage during denoising. To address this, we adopt a restart refinement schedule inspired by diffusion-based sampling methods [xu2023restart]. After several rounds of structure and appearance control, we inject noise at an intermediate timestep $t_{\text{min}}$, effectively restarting the denoising process by transitioning the latent to $t_{\text{max}}$ step. A DDIM backward step [song2020ddim] is then applied. This forward–backward cycle is repeated $N$ times within $t_{\text{min}},t_{\text{max}}]$. In the $i^{\textrm{th}}$ iteration ($i\in\{0,\dots,N-1\}$), the restart proceeds as follows:
@@ 4.3 Appearance-Rich Prompting
To enhance the realism of the generated image, prior work [mo2023freecontrol, lin2024ctrlx] generates an auxiliary appearance image $\mathbf{I}^{\text{app}}$ and perform appearance transfer (the middle branch of the U-Net in Fig. 4; see Appx. D for technical details). However, brief and ambiguous user prompts sometimes hinder the establishment of semantic correspondence between the appearance image $\mathbf{I}^{\text{app}}$ and the output image $\mathbf{I}$ in existing appearance transfer methods, leading to artifacts. For example, as illustrated in Fig. 4, the condition image is a frontal view of a cat’s head, while the text prompt specifies ”a bear”, resulting in an appearance image $\mathbf{I}^{\text{app}}$ that depicts a full-body bear.
To tackle this issue, we propose Appearance-Rich Prompting (ARP), a strategy that leverages multimodal large language models [achiam2023gpt4] to systematically align the semantics of $\mathcal{P}$ with those of the conditions $\mathbf{I}^{\text{struct}}$. Illustrative examples are provided in the ablation study in Fig. 8. Please refer to Appx. D for details of the prompt engineering pipeline $\mathcal{E}$.
@@ 5.1 Setup
We base our evaluation on datasets from prior work [mo2023freecontrol, lin2024ctrlx]. However, many of the condition types in prior datasets are underrepresented, resulting in an imbalanced distribution. To enable more consistent evaluation, we collect seven commonly used structural conditions: canny edges, depth maps, normal maps, human poses, segmentation masks, HED edges, and scribble drawings, and construct a more balanced dataset with over 20 image-text pairs for each condition type.
@@
We evaluate our method against 6 existing training-free baselines: Ctrl-X [lin2024ctrlx], FreeControl [mo2023freecontrol], PnP [tumanyan2023plug-and-play], P2P [hertz2022prompt-to-prompt], SDEdit [meng2021sdedit], and InfEdit [xu2023infedit], as well as 2 training-based baselines: ControlNet [zhang2023controlNet] and T2I-Adapter [mou2023t2i]. Experiment results on four condition types supported by T2I-Adapter-SDXL (canny, depth, normal, and pose) are provided in Sec. F.3.
@@
Following prior work [mo2023freecontrol, lin2024ctrlx], we employ three widely-adopted metrics for comparison. (i) CLIP score [radford2021clip] measures text-image alignment via cosine similarity between the CLIP embeddings of the generated image and text prompt. (ii) DINO self-similarity score (Self-sim) [caron2021dino, tumanyan2022splicing] quantifies structural alignment in the DINO-ViT feature space. (iii) Condition LPIPS score (LPIPS) [zhang2018lpips] measures perceptual deviation between the generated image and the condition image.
In addition to traditional metrics, we further adopt three reward model metrics that more accurately reflect human preferences. (i) DreamSim [fu2023dreamsim] evaluates the perceptual similarity of two images by capturing both mid-level similarities (image layout, object pose, semantic content) and low-level attributes (color, texture). (ii) ImageReward [xu2023imagereward] measures image quality and text-image alignment based on a reward model trained with human feedback. (iii) HPSv2 [wu2023hpsv2] serves as a reliable indicator of overall generation quality aligned with human judgments.